1/1

Caroline Keppel,

Robin Adair

What’s this dull town to me Robin’s not near What was’t I wish’d to see What wish’d to hear

Where all the joy and mirth Made this town heaven on earth Oh, they’re all fled with thee Robin Adair

What made th’ assembly shine Robin Adair What made the ball sae fine Robin was there

What when the play was o’er What made my heart so sore Oh, it was parting with Robin Adair

But now thou’rt cold to me Robin Adair But now thou’rt cold to me Robin Adair

Yet he I loved so well Still in my heart shall dwell Oh, I can ne’er forget Robin Adair

Verkinto de tiu ĉi Angla poemo estas Caroline Keppel (∗1734 – †1769).

Arg-1025-2077 (2014-05-27 15:55:44)

La teksto de tiu ĉi kanto troviĝas en http: // www. traditionalmusic. co. uk/ folk-song-lyrics/ Robin_Adair. htm . Pri Robert Adair vidu la retejon http: // en. wikipedia. org/ wiki/ Robert_Adair_( politician) . Legu ankaŭ: http: // www. lookandlearn. com/ blog/ 24 82/ lady-caroline-keppels-sad-love-song-immortalised-young-robin-adair/ . Vidu ankaŭ: http: // www. bartleby. com/ 360/ 3/ 74. html . La kanto estas aŭdebla en http: // www. youtube. com/ watch? v= 51cI0_exnLU .

http://www.poezio.net/version?poem-id=1025&version-id=2077